The charge relationship for a capacitor is given by the following expression:
1/2QV = 1/2CV²

Find V if Q = 750x10^-6 colombs and C = 6x10^-6 farads
1
Expert's answer
2017-05-02T14:27:09-0400
1/2QV=1/2CV^2
QV=CV^2
Q=CV
V=Q/C=(750*10^(-6))/(6*10^(-6))=750/6=125 (volt)